Three identical isolated capacitors, each with capacitance C, are connected in parallel, with a 11.3 V potential difference across their plates. Each capacitor carries a charge of 4.60 uC. (a) Find C (b) Find the equivalent capacitance. (c) Suppose that one of the capacitors is first disconnected and then reconnected with its terminals reversed. Find the final voltage drop across the capacitors
Answer #1

a)

V = potential difference across each capacitor = 11.3 Volts

Q = charge stored = 4.60 x 10-6 C

Capacitance is given as

C = Q/V = (4.60 x 10-6)/11.3 = 0.41 x 10-6 F

b)

equivalent capacitance is given as

Ceq = C1 + C2 + C3 = C + C + C = 3 C = 3 (0.41 x 10-6) = 1.23 x 10-6

c)

The Voltage drop remains same